OpenStudy (anonymous):

\[\int\limits_{-1}^{5} \frac{ \sqrt{\ln(11-x)} }{ \sqrt{\ln(11-x)} + \sqrt{\ln(x+7)}}dx\]

hartnn (hartnn):

interesting.

hartnn (hartnn):

use, \(\large \int \limits_a^b f(x)dx =\int \limits_a^b f(a+b-x)dx \)

hartnn (hartnn):

then add.

hartnn (hartnn):

I mean write your original definite integral as I=....... -->(1) then apply that property, to get I= ... --->(2) then ADD (1) and (2)

hartnn (hartnn):

if i am not wrong, you'll get \(2I=\int \limits_{-1}^51.dx\) and then it'll be easy to find I.

OpenStudy (anonymous):

\[f(x) = \sqrt{\ln(11-x)} , f(4-x) = \sqrt{\ln(x+7}\] and those are defined at the [-1,5] range?

hartnn (hartnn):

no, f(x) is not sqrt{ln(11-x)} f(x) will be entire function, including numerator and denom...

hartnn (hartnn):

\(\large f(x)=\frac{ \sqrt{\ln(11-x)} }{ \sqrt{\ln(11-x)} + \sqrt{\ln(x+7)}}\)

hartnn (hartnn):

what is f(a+b-x) =f(4-x) =.. ?

OpenStudy (anonymous):

hmm seems simplier to do it your way. I thought if we could write it that way then I could get into a position like that : int(-1,0) f(x) / (f(x) + (f(4-x) )dx + int(0,4) f(x)/(f(x)+ f(4-x))dx + int(4,5)f(x)/(f(x)+ f(4-x))dx as those functions are contionues at the -1,5 , taking them into some phrantesis and doing some process we can maybe have this : \[\int\limits_{-1}^{0}\frac{ f(x) + f(4-x) }{ f(x) + f(4-x) }dx + \int\limits_{0}^{2}\frac{ f(x) + f(4-x) }{ f(x) + f(4-x) dx }\] which equals to : \[\int\limits_{-1}^{0}dx + \int\limits_{0}^{2}dx = 1 + 2 = 3\] ain't sure if it's corret or can be done

OpenStudy (anonymous):

(messed around wit du's also there)

hartnn (hartnn):

how does your numerator contain both f(x)+f(4-x) ? u got my method ?

OpenStudy (anonymous):

yeah I understoodyour method

OpenStudy (anonymous):

@hartnn is this a formula?